# A mesh-free hybrid Chebyshev-Tucker tensor format with applications to multi-particle modelling

### v1

@@ 4.2.2 $O(\log N)$ scaling of the Tucker rank of the ChebTuck format
In this section, we demonstrate numerically that the Tucker rank of our mesh-free ChebTuck format of the long-range part of the total potential in a bio-molecular system scales logarithmically with the number of particles $N$, consistent with the result for grid-based range-separated format as in [4, Theorem 3.1]. We consider a bio-molecular system in $d=3$ dimension with $N$ particles discretized on Cartesian grid with grid size $n=2048$. For varying particle number $N=100,\cdots,600$ and Chebyshev degree $m=3,5,9,\cdots,1025$, we show the Tucker rank of the ChebTuck format with $\varepsilon=10^{-7}$ RHOSVD truncation tolerance in Table 4.6. We can see that the Tucker rank grows only mildly with respect to the number of particles $N$. To have a closer look at the actual behavior, we also plot the singular values of the mode-$2$ side matrices of the Chebyshev coefficient tensor ${\bf C}$ (which is in CP format by construction) for $m=129$ in Fig. 4.9. We can see that the singular values drop to the level of machine precision after the first few significant ones, which also explains the behavior of Fig. 4.6, middle figure.
$N$ $m$ 9 17 33 65 129 257 513 1025
@@
Table 4.6: Tucker rank $\max_{\ell=1,2,3}{r_{i_{\ell}}}$ of the ChebTuck format with RHOSVD truncation tolerance $10^{-7}$ for the long-range part of the total potential in a bio-molecular with $n=2048$ for different particle number $N$ and Chebyshev degree $m$.
Figure 4.9: Left: Singular values of the mode-2 side matrices of the Chebyshev coefficient tensor ${\bf C}\in\mathbb{R}^{m\times m\times m}$ for $m=129$. Right: the zoom-in from the 10th to the 40th singular values of the same plot.
@@ 4.2.3 A lattice-type compound with vacancies
In this section, we apply our hybrid tensor format ChebTuck to the problem of low-parametric approximation of electrostatic potentials of the large lattice-type structures, see [25]. In this case, the charge centers $x_{v}$ in Eq. 4.1 are located on a regular grid. For a lattice-type structure of size $24\times 24\times 4$ with $N=2304$ particles, the canonical representation of its potential is given by a CP tensor of size $n\times n\times n$ with $n=960$ and rank $R=50$.
We apply Algorithm 4 to approximate the potential with Chebyshev degrees $m_{1}=m_{2}=m_{3}=m=129$ and Tucker-ALS truncation tolerance $10^{-7}$. The results are shown in Fig. 4.10. We observe that the error of our method is of the magnitude $10^{-3}$, which is not as good as the bio-molecule example since we have not separated the long-range and short-range parts. On the other hand, see Fig. 4.11, if we separate the long-range part of the potential by taking out the last $22$ canonical vectors, the error is reduced to the magnitude $10^{-6}$, which is consistent with the prescribed tolerance and previous results on the bio-molecule example.
The numerical results confirm the efficiency of the new mesh-less hybrid tensor decomposition on the example of complicated multi-particle potential arising in quantum chemistry, bio-molecular modeling, etc. Indeed, it demonstrates the same approximation error as for the established grid-based tensor approximation method applied on relatively large grids, required for the good resolution of the complicated multi-particle potentials.
Figure 4.10: Left: the middle slice $\hat{f}_{\bf m}(:,:,t_{n/2})\approx{\bf A}(:,:,n/2)$ of the degree $m=129$ ChebTuck format for the lattice structure with $N=2304$ and $n=960$ without range separation. Middle: the error during RHOSVD compression, i.e. $\hat{f}_{\bf m}(:,:,t_{n/2})-\tilde{f}_{\bf m}(:,:,t_{n/2})$. Right: the total error $\hat{f}_{\bf m}(:,:,t_{n/2})-{\bf A}_{R_{l}}(:,:,n/2)$.
Figure 4.11: Left: the middle slice $\hat{f}_{\bf m}(:,:,t_{n/2})\approx{\bf A}_{R_{l}}(:,:,n/2)$ of the degree $m=129$ ChebTuck format for the long-range part of a lattice structure with $N=2304$ and $n=960$. Middle: the error during RHOSVD compression, i.e. $\hat{f}_{\bf m}(:,:,t_{n/2})-\tilde{f}_{\bf m}(:,:,t_{n/2})$. Right: the total error $\hat{f}_{\bf m}(:,:,t_{n/2})-{\bf A}_{R_{l}}(:,:,n/2)$.
@@ 5 Conclusions
We introduce and analyze a mesh-free two-level hybrid Tucker tensor format for approximation of multivariate functions in $\mathbb{R}^{3}$ with both high and reduced regularity, which combines the tensor product Chebyshev polynomial interpolation in the global computational domain with the ALS-based Tucker decomposition of the Chebyshev coefficient tensor.
For functions with multiple cusps (or even stronger singularities) our method applies to the long-range component of the target function represented in low-rank RS tensor format. In case of orthogonal set of interpolating functions like Chebyshev polynomials, the subsequent rank optimization via ALS-type rank-structured decomposition of the coefficient tensor leads to nearly optimal $\varepsilon$-rank Tucker decomposition of the initial function with merely the same rank parameter as for the alternative grid-based Tucker-ALS approximation.
We present numerous numerical tests demonstrating the efficiency of the proposed techniques on the examples of many-particle electrostatic potentials in bio-molecular systems and in lattice-structured compounds. In particular, we observe from numerical examples the following beneficial features of our method.
  • Fast mesh-free Tucker decomposition with quasi-optimal rank parameter in application to rather regular functions.
  • Efficient mesh-free Tucker tensor decomposition of the long-range part for the multi-particle electrostatic potentials in $\mathbb{R}^{3}$, composed of the large sum of singular Newton kernels.
  • The new mesh-free two-level hybrid Tucker tensor format demonstrates the efficiency in application to lattice-structured compounds with monopoles and dipoles like structures.
  • In case of input function presented in the canonical (CP) tensor format our method needs only a small number of 1D-Chebyshev interpolations of the 1D-canonical skeleton vectors.
Finally, we notice that the presented techniques can be applied to the Green functions for various elliptic operators and can be extended to higher dimensions. In the latter case, the combination with the TT, QTT or H-Tucker tensor decompositions is necessary to avoid the curse of dimensionality.

### v2

@@ 4.2.2 $O(\log N)$ scaling of the Tucker rank of the ChebTuck format
In this section, we demonstrate numerically that the Tucker rank of our mesh-free ChebTuck format of the long-range part of the total potential in the biomolecule scales logarithmically with the number of particles $N$, consistent with discussions in Section 3.2.3 as well as the result for grid-based range-separated format as in [6, Theorem 3.1]. We consider the biomolecule in $d=3$ dimension with $N$ particles discretized on Cartesian grid with grid size $n=2048$. For varying particle number $N=100,\ldots,600$ and Chebyshev degree $m=3,5,9,\ldots,1025$, we show the Tucker rank of the ChebTuck format with $\varepsilon=10^{-7}$ RHOSVD truncation tolerance in Table 4. We can see that the Tucker rank grows only mildly with respect to the number of particles $N$. To have a closer look at the actual behavior, we also plot the singular values of the mode-$2$ side matrices of the Chebyshev coefficient tensor $\bf C$} (which is in CP format by construction) for $m=129$ in Fig. 8. We can see that the singular values drop to the level of machine precision after the first few significant ones, which also explains the behavior of the middle plot in Fig. 5.
NN $m$ 9 17 33 65 129 257 513 1025
@@
Table 4: Tucker rank $\max_{\ell=1,2,3}{r_{i_{\ell}}}$ of the ChebTuck format with RHOSVD truncation tolerance $10^{-7}$ for the long-range part of the total potential in a bio-molecule with $n=2048$ for different particle numbers NN and Chebyshev degrees $m$.
Figure 8: Left: Singular values of the mode-2 side matrices of the Chebyshev coefficient tensor ${\bf C}\in\mathbb{R}^{m\times m\times m}$ with $m=129$ for the long-range part of the total potential in the biomolecule with $n=2048$. Right: the zoom-in from the 10th to the 40th singular values of the same plot.
@@ 4.2.3 A lattice-type compound with vacancies
Finally, we apply our hybrid tensor format ChebTuck to the problem of low-parametric approximation of electrostatic potentials of large lattice-type structures, see [33] for tensor-based approach to lattice summation of electrostatic potentials. Fast computation of lattice potentials was since longer an important topic of numerical analysis because of various applications in crystal modeling, in quantum chemistry computations, in polymer simulations and many other fields. Traditional numerical methods for lattice summation are based on classical fast Ewald summation method via transition of the problem to the Fourier space [23, 14, 31], or, alternatively, by using the fast multipole method [26].
Here we demonstrate our new approach in application to the long-range part of the many-particles lattice-structured electrostatic potential. In this case, the charge centers $x_{v}$ in Eq. 33 are located on a regular grid. For a lattice-type structure of size $24\times 24\times 4$ with $N=2304$ particles, the canonical representation of its potential is given by a CP tensor of size $n\times n\times n$ with $n=960$ and rank $R=50$. We apply Algorithm 3 to approximate the long-range part of the potential with Chebyshev degrees $m_{1}=m_{2}=m_{3}=m=129$ and Tucker-ALS truncation tolerance $10^{-7}$. The results are shown in Fig. 9. We observe that the error is of magnitude $10^{-6}$, which is consistent with the prescribed tolerance and the previous results on the bio-molecule example. The numerical results confirm the efficiency of the new mesh-less hybrid tensor decomposition on the example of complicated multi-particle potential arising in quantum chemistry, biomolecule modeling, etc. Indeed, it demonstrates the same approximation error as for the established grid-based tensor approximation method applied on relatively large grids, required for the good resolution of complicated multi-particle potentials.
Figure 9: Left: the middle slice $\hat{f}_{\bf m}(:,:,t_{n/2})\approx{\bf A}_{R_{l}}(:,:,n/2)$ of the degree $m=129$ ChebTuck format for the long-range part of a lattice structure with $N=2304$ and $n=960$. Middle: the error during RHOSVD compression, i.e. $\hat{f}_{\bf m}(:,:,t_{n/2})-\tilde{f}_{\bf m}(:,:,t_{n/2})$. Right: the total error $\hat{f}_{\bf m}(:,:,t_{n/2})-{\bf A}_{R_{l}}(:,:,n/2)$.
@@ 5 Conclusions
We introduce and analyze a mesh-free two-level hybrid Tucker tensor format for approximation of trivariate functions on a hypercube with both high and reduced regularity, which combines the tensor product Chebyshev polynomial interpolation in the global computational domain with the ALS-based Tucker decomposition of the Chebyshev coefficient tensor. For functions with multiple cusps (or even stronger singularities), our method applies to the long-range component of the target function represented in low-rank RS tensor format. The Tucker rank optimization via ALS-type rank-structured decomposition of the coefficient tensor leads to nearly optimal $\varepsilon$-rank Tucker decomposition of the initial function with merely the same rank parameter as for the alternative grid-based methods. We present numerous numerical tests demonstrating the efficiency of the proposed techniques on the examples of many-particle electrostatic potentials in biomolecular systems and in lattice-structured compounds. In case of input functions presented in the canonical (CP) tensor format, our method needs only 1D-Chebyshev interpolations of the 1D-canonical vectors.
Finally, we notice that the presented techniques can be applied to the Green functions for various elliptic operators and can be extended to higher dimensions. In the latter case, the combination with cross-based tensor decompositions in TT, QTT or H-Tucker formats is necessary to avoid the curse of dimensionality.
Reproducibility statement
The full source code and data to reproduce the numerical experiments is available at https://github.com/bonans/ChebTuck.